OpenStudy (anonymous):

which expression represents 8x^3/7 in radical form

OpenStudy (jdoe0001):

$$ \large 8^{\frac{3}{7}} ? $$

OpenStudy (anonymous):

also What is the equation of a horizontal line passing through (–5, 3)? x = –5 y = –5 x = 3 y = 3

OpenStudy (anonymous):

yes @jdoe0001

OpenStudy (jdoe0001):

$$ \large 8^{\frac{\color{blue}{n}}{\color{red}{m}}} \implies \sqrt[\color{red}{m}]{8^{\color{blue}{n}}} $$

OpenStudy (jdoe0001):

so, what is it?

OpenStudy (anonymous):

\[8^{\frac{ 3 }{7 }}\]

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@tcarroll010

OpenStudy (jdoe0001):

well, that's what you have, what's required is the "radical" form, not the "exponential" form

OpenStudy (anonymous):

@jdoe0001 how do you get that colored text?

OpenStudy (jdoe0001):

\large 8^{\frac{\color{blue}{n}}{\color{red}{m}}} \implies \sqrt[\color{red}{m}]{8^{\color{blue}{n}}} :) is part of the library used here :), I mainly use red/green/blue, but it takes #aabbcc hex values too \color {#a020f0}{25} or \(\huge \color {#a020f0}{25} \)

OpenStudy (jdoe0001):

@FutureMathProfessor